was Dingbat 101 \dingfill (with the usual numeric parameter) acts like other filling com-

mands, but fills the space with a chosen Dingbat like that.

\dingline generates a freestanding line of the little chaps:

Inevitably, the ‘dinglist’ environment sets up an itemized list, but has a Dingbat instead of

a bullet (you are reading a dinglist at the moment). ϑ ε was symbol 101 ϑ \Pifill{psy} (with the usual numeric parameter) acts like other filling

commands, but fills the space with a chosen symbol ε ε ε like that. ϑ \Piline generates a freestanding line of the little chaps:

, , , , , , , , , , , , , , , ,

ϑ Inevitably, the ‘Pilist’ environment sets up an itemized list, but has a

symbol instead of a bullet (you are reading a Pilist at the moment). text inside list,

α text inside list, β text inside list,

i. text inside list, text inside list, text inside list,

ii. text inside list, χ text inside list,

text inside list So we can refer to item as usual.

\Pisymbol generates a Pi character; it has two parameters: the font family, and the character number. Thus \Pisymbol{pzd}{'166} generates .

\Pifill (with the same parameters) acts like the other filling commands

in TEX, but fills the space with a chosen symbol like that. \Piline generates a freestanding line of the little chaps: